# ELemental abundances of Planets and brown dwarfs Imaged around Stars (ELPIS): II. The Jupiter-like Inhomogeneous Atmosphere of the First Directly Imaged Planetary-Mass Companion 2MASS 1207 b

### v1

@@ Abstract
2MASS 1207 b, the first directly imaged planetary-mass companion, has been instrumental in advancing our understanding of exoplanets and brown dwarfs over the past 20 years. We have performed extensive atmospheric retrieval analyses of 2MASS 1207 b’s JWST/NIRSpec spectrum using petitRADTRANS and a new atmospheric inhomogeneity framework, which characterizes homogeneous atmospheres, patchy clouds, cloud-free hot spots, or the combination of patchy clouds and spots. Among 24 retrieval runs with various assumptions, the most statistically preferred model corresponds to the patchy cloud scheme, with $T_{\rm eff}=1174^{+4}_{-3}$ K, $\log{(g)}=3.62^{+0.03}_{-0.02}$ dex, and $R=1.399^{+0.008}_{-0.010}$ RJup, along with near-solar atmospheric compositions of [M/H]$=-0.05\pm 0.03$ dex and C/O$=0.440\pm 0.012$. This model suggests $\sim 9\%$ of 2MASS 1207 b’s atmosphere is covered by thin iron clouds, producing L-dwarf-like spectra, while the remaining $91\%$ consists of thick silicate and iron clouds, emitting blackbody-like spectra. These thin-cloud patches and thick-cloud regions resemble Jupiter’s belts and zones, respectively, and this scenario is consistently supported by other retrieval runs incorporating inhomogeneous atmospheres. We demonstrate that the weak CO absorption of 2MASS 1207 b can be explained by the veiling effects of patchy thick clouds; the absence of 3.3 $\mu$m CH4 absorption is attributed to its hot thermal structure, which naturally leads to a CO-dominant, CH4-deficient atmosphere. The retrieved atmospheric models also match the observed variability amplitudes of 2MASS 1207 b. Our analysis reveals that the inferred atmospheric properties show significant scatter in less statistically preferred retrieval runs but converge to consistent values among the preferred ones. This underscores the importance of exploring diverse assumptions in retrievals to avoid biased interpretations of atmospheric properties and formation pathways.
journal: AAS Journalsfacilities: JWST (NIRSpec)software: petitRADTRANS (Mollière et al., 2019; Nasedkin et al., 2024a), easyCHEM (Mollière et al., 2017; Lei & Mollière, 2024), MultiNest (Feroz & Hobson, 2008; Feroz et al., 2009, 2019), PyMultiNest (Buchner et al., 2014), corner.py (Foreman-Mackey, 2016), astropy (Astropy Collaboration et al., 2013, 2018), ipython (Pérez & Granger, 2007), numpy (Oliphant, 2006), scipy (Jones et al., 2001), matplotlib (Hunter, 2007).
@@ 2 A Brief Review of 2MASS 1207 A+b
The planetary-mass companion, 2MASS 1207 b (L6), was first discovered by Chauvin et al. (2004) through high-contrast adaptive optics imaging, with its physical association to 2MASS 1207 A confirmed by Chauvin et al. (2005) and Song et al. (2006). As the first directly imaged planetary-mass companion, 2MASS 1207 b has received numerous follow-up observations focusing on its thermal emission spectrophotometry (e.g., Chauvin et al., 2004; Song et al., 2006; Mohanty et al., 2007; Patience et al., 2010; Skemer et al., 2011, 2014; Zhou et al., 2016, 2019; Luhman et al., 2023; Manjavacas et al., 2024).
@@
Accretion emission lines, such as He i and Paschen series, were recently revealed in 2MASS 1207 b’s JWST/NIRSpec spectrum (Luhman et al., 2023; Manjavacas et al., 2024; Marleau et al., 2024), indicating the presence of a circum-secondary disk. The W2 photometry (4.6 $\mu$m) of 2MASS 1207 b does not show strong excess, suggesting that the disk emission contributes to its SED beyond 5 $\mu$m (Luhman et al., 2023). The mass of 2MASS 1207 b’s dust disk is likely below 1 Mmoon based on ALMA observations Ricci et al. (2017).
@@ 2.3 The 2MASS 1207 A+b Ecosystem
The mass ratio between 2MASS 1207 b and 2MASS 1207 A ($M_{b}/M_{A}\approx 0.2$; e.g., Chauvin et al., 2005) suggests that this system is more analogous to stellar binaries than to typical planetary systems discovered by transits or radial velocity methods (see also Figure 1 of Zhang et al., 2025). Any additional companions with masses $>2$ $M_{\rm Jup}$ and orbital separation $>75$ au were ruled out by Zhou et al. (2019). For 2MASS 1207 b to have formed via disk instability, an unusually high initial disk mass for 2MASS 1207 A would be required (e.g., Gizis et al., 2005). Furthermore, the core accretion mechanism would likely be inefficient in assembling 2MASS 1207 b’s mass at its current orbit ($\approx 50$ au) within its young age (e.g., Lodato et al., 2005). However, outward migration could alleviate some of the tension regarding the formation timescale (e.g., Martin et al., 2007). Consequently, it is widely theorized that 2MASS 1207 A+b formed via gravitational instability during the collapse of a molecular cloud (e.g., Lodato et al., 2005; Gizis et al., 2005; Chauvin et al., 2005; Boss, 2006; Stamatellos & Whitworth, 2009). This scenario is further supported by the fact that the mass of 2MASS 1207 b is slightly above the estimated lowest mass threshold for star formation products (e.g., Zuckerman & Song, 2009; Kirkpatrick et al., 2021).
@@ 3 Archival JWST/NIRSpec Spectroscopy of 2MASS 1207 b
For the subsequent retrieval analysis, we masked a few accretion emission lines with detection significance values exceeding $4\sigma$ (Figure 1), as identified by Marleau et al. (2024). This includes Pa $\alpha$ (1.874–1.8765 $\mu$m), Pa $\beta$ (1.2812–1.2836 $\mu$m), Pa $\gamma$ (1.0935–1.0948 $\mu$m), Pa $\delta$ (1.004–1.0058 $\mu$m), and the He i triplet (1.0824–1.0842 $\mu$m). Additionally, 2MASS 1207 b does not exhibit excess emission in the W2 band (4.6 $\mu$m), suggesting that its potential circumstellar disk does not significantly contribute to the JWST/NIRSpec spectroscopy (e.g., Luhman et al. 2023; see also Section 2.2).
@@ 5.2.2 An Expanded Grid of Chemical Equilibrium Abundances
In this work, we expand the petitRADTRANS chemical equilibrium abundance grid over an even broader parameter space and include more opacity sources. This new grid is computed using easyCHEM [^3] https://easychem.readthedocs.io/en/latest/ (Mollière et al., 2017; Lei & Mollière, 2024). Our grid spans [40 K, 6000 K] in temperature ($T$), [$10^{-8}$ bar, $10^{3}$ bar] in pressure ($P$), [$-2$ dex, $+3$ dex] in [M/H], and [$0.1$, $1.6$] in C/O, with 120, 100, 40, and 20 grid points, respectively. These grid points are evenly spaced logarithmically for $T$ and $P$ and linearly for [M/H] and C/O. Compared to the grid from Mollière et al. (2019), this new grid extends to higher temperatures and includes additional species, such as AlH, C2, CN, OH, Al, Ca, Fe, Mg, Si, and Ti.
@@ 5.3 Cloud Model
Our retrievals are performed using both cloud-free and cloudy scenarios. For the cloudy setups, we test various combinations of cloud species, including amorphous silicates [^4] We focus on Mg2SiO4 and MgSiO3 clouds in this work but note that the upcoming MIRI spectra will better constrain the specific type of silicate clouds in 2MASS 1207 b’s atmosphere (GTO program 1270). of forsterite (Mg2SiO4) and enstatite (MgSiO3) stoichiometry (Jäger et al., 2003) and iron (Fe; Pollack et al., 1994) with Mie scattering particle shapes. Additionally, in some cloudy retrieval runs, we incorporate a gray cloud layer with a varying pressure $P_{\rm gray}$; the opacity for deeper pressure layers with $P>P_{\rm gray}$, $\kappa_{\rm gray}$, is introduced as a free parameter, assumed to remain constant as a function of pressures.
@@ 4. $\bullet$
In our analysis, the lowest BIC corresponds to a patchy cloud scenario with inhomogeneous forsterite and iron clouds, identified as the scheme (b) in our atmospheric inhomogeneity framework (Section 5.4). This run outperforms other runs with $\Delta$BIC values ranging from 4 to nearly 5000, making it the most favored model for interpreting the JWST NIRSpec spectrum of 2MASS 1207 b. Throughout this manuscript, we refer to this most preferred model as QEQ-1 and label other, less favored models as QEQ-x, where x denotes their ranks by BIC values. A larger x corresponds to a higher $\Delta$BIC and a less preferred model. We compare the observations with the fitted model spectra of QEQ-1 in Figure 6 and present the posterior distributions of key retrieved parameters in Figure 7.
@@ 2. $\bullet$
The thick clouds and potential high-altitude gray cloud content of 2MASS 1207 b, as suggested by our retrievals, may be contributed by this object’s ongoing accretion (Luhman et al., 2023; Marleau et al., 2024; Manjavacas et al., 2024) of dust and gas from the circumstellar disk of 2MASS 1207 A (e.g., Ricci et al., 2017) which has an estimated dust mass of $\sim 0.1$ M (e.g., Mohanty et al., 2013; Ricci et al., 2017). Additionally, 2MASS 1207 b may host its own disk, though if present, it likely has a low dust mass below 1 Mmoon (Ricci et al., 2017). Similar to events that have influenced Jupiter — such as the impacts of Comet Shoemaker-Levy 9 (e.g., Hammel et al., 1995) — such processes may also influence the atmospheric evolution of 2MASS 1207 b. Follow-up studies similar to Madurowicz et al. (2023) would be valuable in assessing whether accretion activity can match the observed cloud properties of 2MASS 1207 b, as well as the present-day atmospheric compositions of both 2MASS 1207 A and b. Furthermore, future retrieval studies that incorporate both NIRSpec and MIRI spectra of 2MASS 1207 b will be crucial for more precisely constraining the type and amount of silicate clouds in its atmosphere.
@@ 7.2.2 Absent CH4 absorption due to the hot atmospheric thermal structure
In Figure 10, we compare the retrieved T-P profile with a critical curve corresponding to an equilibrium situation where the volume mixing ratios of CH4 and CO are equal. This curve, $\log{(P)}=5.05-5807.5/T+0.5\times$[M/H], was derived by Fortney et al. (2020). Deviations from this curve lead to either a decrease in CH4 abundance at higher temperatures and lower pressures, or a decrease in CO abundance at lower temperatures and higher pressures. Objects with thermal structures hotter than this CO–CH4 transition curve tend to exhibit weak or absent 3.3 $\mu$m CH4 absorption, while cooler objects display stronger CH4 absorption. We demonstrate this effect in Figures 1112 using the Sonroa Diamondback atmospheric models (Morley et al., 2024).
@@ 7.5 Atmospheric Composition
Focusing on retrieval runs with the seven lowest BIC values (QEQ-1 to QEQ-7), the inferred [M/H] of 2MASS 1207 b ranges from $-0.12$ to $+0.10$ dex, and C/O from 0.44 to 0.52 (Table3). These retrievals incorporate models of atmospheric inhomogeneity, such as patchy clouds or a combination of patchy clouds and hot spots (Figure 5), and explore various assumptions about cloud composition and parameter priors. Despite these differences in model assumptions, their consistent results support the robustness of the near-solar [M/H] and C/O values. Specifically, the most favored retrieval setup, QEQ-1, yields a [M/H]$=-0.05\pm 0.03$ dex and a C/O of 0.440±0.012plus-or-minus0.4400.0120.440$\pm$ 0.0120.440 ± 0.012.888The inferred atmospheric compositions of 2MASS 1207 b — and more broadly, those of exoplanets and brown dwarfs — are derived from disk-integrated spectra, which average out potential three-dimensional spatial variations in atmospheric abundances. Studies of Jupiter suggested that certain gas abundances retrieved from disk-integrated spectra can differ from those of localized regions (e.g., Gapp et al., 2024; Irwin et al., 2025). Comparing the globally and locally inferred atmospheric properties of solar-system planets, such as Jupiter, provides crucial context for interpreting exoplanet and brown dwarf atmospheres and remains a key topic of ongoing research. As a member of the TW Hydrae association, 2MASS 1207 b’s near-solar [M/H] is consistent with stellar members of other nearby young moving groups (e.g., Santos et al., 2008; Ammler-von Eiff & Guenther, 2009; Viana Almeida et al., 2009; D’Orazi et al., 2011; Biazzo et al., 2011, 2012a, 2017).

### v2

@@ Abstract
2MASS 1207 b, the first directly imaged planetary-mass companion, has been instrumental in advancing our understanding of exoplanets and brown dwarfs over the past 20 years. We have performed extensive atmospheric retrieval analyses of 2MASS 1207 b’s JWST/NIRSpec spectrum using petitRADTRANS and a new atmospheric inhomogeneity framework, which characterizes homogeneous atmospheres, patchy clouds, cloud-free hot spots, or the combination of patchy clouds and spots. Among 24 retrieval runs with various assumptions, the most statistically preferred model corresponds to the patchy cloud scheme, with $T_{\rm eff}=1174^{+4}_{-3}$ K, $\log{(g)}=3.62^{+0.03}_{-0.02}$ dex, and $R=1.399^{+0.008}_{-0.010}$ RJup, along with near-solar atmospheric compositions of [M/H]$=-0.05\pm 0.03$ dex and C/O$=0.440\pm 0.012$. This model suggests $\sim 9\%$ of 2MASS 1207 b’s atmosphere is covered by thin iron and silicate clouds, producing L-dwarf-like spectra, while the remaining $91\%$ consists of thick iron and silicate clouds, emitting blackbody-like spectra. These thin-cloud patches and thick-cloud regions resemble Jupiter’s belts and zones, respectively, and this scenario is consistently supported by other retrieval runs incorporating inhomogeneous atmospheres. We demonstrate that the weak CO absorption of 2MASS 1207 b can be explained by the veiling effects of patchy thick clouds; the absence of 3.3 $\mu$m CH4 absorption is attributed to its hot thermal structure, which naturally leads to a CO-dominant, CH4-deficient atmosphere. The retrieved atmospheric models also match the observed variability amplitudes of 2MASS 1207 b. Our analysis reveals that the inferred atmospheric properties show significant scatter in less statistically preferred retrieval runs but converge to consistent values among the preferred ones. This underscores the importance of exploring diverse assumptions in retrievals to avoid biased interpretations of atmospheric properties and formation pathways.
journal: AJ (Accepted: May 31, 2025)facilities: JWST (NIRSpec)software: petitRADTRANS (Mollière et al., 2019; Nasedkin et al., 2024a), easyCHEM (Mollière et al., 2017; Lei & Mollière, 2024), MultiNest (Feroz & Hobson, 2008; Feroz et al., 2009, 2019), PyMultiNest (Buchner et al., 2014), corner.py (Foreman-Mackey, 2016), astropy (Astropy Collaboration et al., 2013, 2018), ipython (Pérez & Granger, 2007), numpy (Oliphant, 2006), scipy (Jones et al., 2001), matplotlib (Hunter, 2007).
@@ 2 A Brief Review of 2MASS 1207 A+b
The planetary-mass companion, 2MASS 1207 b (L6), was first discovered by Chauvin et al. (2004) through high-contrast adaptive optics imaging, with its physical association to 2MASS 1207 A confirmed by Chauvin et al. (2005) and Song et al. (2006). Based on evolution models and 2MASS 1207 b’s near-infrared absolute magnitudes, Chauvin et al. (2004) and Chauvin et al. (2005) estimated that this object has an effective temperature of $T_{\rm eff}=1200\pm 200$ K, a mass of $M=5\pm 2$ MJup, a bolometric luminosity of $\log{(L/L_{\odot})}\approx-4.28$ dex, and a radius of $R\approx 1.5$ RJup. As the first directly imaged planetary-mass companion, 2MASS 1207 b has received numerous follow-up observations focusing on its thermal emission spectrophotometry (e.g., Chauvin et al., 2004; Song et al., 2006; Mohanty et al., 2007; Patience et al., 2010; Skemer et al., 2011, 2014; Zhou et al., 2016, 2019; Luhman et al., 2023; Manjavacas et al., 2024).
@@
Accretion emission lines, such as He i and Paschen series, were recently revealed in 2MASS 1207 b’s JWST/NIRSpec spectrum (Luhman et al., 2023; Manjavacas et al., 2024; Marleau et al., 2024). The strengths of these lines suggest a mass accretion rate of several ×$\times 10^{-9}$ MJup yr-1, with an estimated uncertainty of about an order of magnitude (Marleau et al., 2024; Aoyama et al., 2024). These emission lines are suggested to originate from magnetospheric accretion rather than from active accretion through a potential circumplanetary disk (Marleau et al., 2024; Aoyama et al., 2024). The presence of such a disk remains unconfirmed at the time of writing. Indeed, the W2 photometry (4.6 $\mu$m) of 2MASS 1207 b does not show strong excess, suggesting that the disk emission, if any, contributes to its SED beyond 5 $\mu$m (Luhman et al., 2023). The mass of 2MASS 1207 b’s dust disk is likely below 1 Mmoon based on ALMA observations (Ricci et al., 2017).
@@ 2.3 The 2MASS 1207 A+b Ecosystem
The mass ratio between 2MASS 1207 b and 2MASS 1207 A ($M_{b}/M_{A}\approx 0.2$; e.g., Chauvin et al., 2005) suggests that this system is more analogous to stellar binaries than to typical planetary systems discovered by transits or radial velocity methods (see also Figure 1 of Zhang et al., 2025). Any additional companions with masses $>2$ $M_{\rm Jup}$ and orbital separation $>75$ au were ruled out by Zhou et al. (2019). For 2MASS 1207 b to have formed via disk instability, an unusually high initial disk mass for 2MASS 1207 A would be required (e.g., Gizis et al., 2005). Furthermore, the core accretion mechanism would likely be inefficient in assembling 2MASS 1207 b’s mass at its current orbit ($\approx 50$ au) within its young age (e.g., Lodato et al., 2005). However, outward migration could alleviate some of the tension regarding the formation timescale (e.g., Martin et al., 2007). Consequently, it is widely theorized that 2MASS 1207 A+b formed via gravitational instability during the collapse of a molecular cloud (e.g., Lodato et al., 2005; Gizis et al., 2005; Chauvin et al., 2005; Boss, 2006; Stamatellos & Whitworth, 2009). This scenario is further supported by the fact that the mass of 2MASS 1207 b is close to the estimated lowest mass threshold for star formation products (e.g., Zuckerman & Song, 2009; Kirkpatrick et al., 2021; Miret-Roig et al., 2022; Pearson & McCaughrean, 2023; Bouy et al., 2025).
@@ 3 Archival JWST/NIRSpec Spectroscopy of 2MASS 1207 b
For the subsequent retrieval analysis, we masked a few accretion emission lines with detection significance values exceeding $3.5\sigma$ (Figure 1), as identified by Marleau et al. (2024). This includes Pa $\alpha$ (1.874–1.8765 $\mu$m), Pa $\beta$ (1.2812–1.2836 $\mu$m), Pa $\gamma$ (1.0935–1.0948 $\mu$m), Pa $\delta$ (1.004–1.0058 $\mu$m), and the He i triplet (1.0824–1.0842 $\mu$m). Additionally, 2MASS 1207 b does not exhibit excess emission in the W2 band (4.6 $\mu$m), suggesting that its potential circumstellar disk does not significantly contribute to the JWST/NIRSpec spectroscopy (e.g., Luhman et al. 2023; see also Section 2.2).
@@ 5.2.2 An Expanded Grid of Chemical Equilibrium Abundances
In this work, we expand the petitRADTRANS chemical equilibrium abundance grid over an even broader parameter space and include more opacity sources. This new grid is computed using easyCHEM [^3] https://easychem.readthedocs.io/en/latest/ (Mollière et al., 2017; Lei & Mollière, 2024). Our grid spans [40 K, 6000 K] in temperature ($T$), [$10^{-8}$ bar, $10^{3}$ bar] in pressure ($P$), [$-2$ dex, $+3$ dex] in [M/H], and [$0.1$, $1.6$] in C/O, with 120, 100, 40, and 20 grid points, respectively. These grid points are evenly spaced logarithmically for $T$ and $P$ and linearly for [M/H] and C/O. Compared to the grid from Mollière et al. (2019), this new grid extends to higher temperatures and includes additional species, such as AlH, C2, CN, OH, Al, Ca, Fe, Mg, Si, and Ti. This grid is available via the Zenodo repository, https://doi.org/10.5281/zenodo.15654694.
@@ 5.3 Cloud Model
Our retrievals are performed using both cloud-free and cloudy scenarios. For the cloudy setups, we test various combinations of cloud species, including amorphous silicates [^4] We focus on Mg2SiO4 and MgSiO3 clouds in this work but note that the upcoming MIRI spectra will better constrain the specific type of silicate clouds in 2MASS 1207 b’s atmosphere (GTO program 1270), as well as its potential circumsecondary disk. of forsterite (Mg2SiO4) and enstatite (MgSiO3) stoichiometry (Jäger et al., 2003) and iron (Fe; Pollack et al., 1994) with Mie scattering particle shapes. Additionally, in some cloudy retrieval runs, we incorporate a gray cloud layer with a varying pressure $P_{\rm gray}$; the opacity for deeper pressure layers with $P>P_{\rm gray}$, $\kappa_{\rm gray}$, is introduced as a free parameter, assumed to remain constant as a function of pressures.
@@ 4. $\bullet$
In our analysis, the lowest BIC corresponds to a patchy cloud scenario with inhomogeneous forsterite and iron clouds, identified as the scheme (b) in our atmospheric inhomogeneity framework (Section 5.4). This run outperforms other runs with $\Delta$BIC values ranging from 4 to nearly 5000, making it the most favored model for interpreting the JWST NIRSpec spectrum of 2MASS 1207 b. Throughout this manuscript, we refer to this most preferred model as QEQ-1 and label other, less favored models as QEQ-x, where x denotes their ranks by BIC values. A larger x corresponds to a higher $\Delta$BIC and a less preferred model. We compare the observations with the fitted model spectra of QEQ-1 in Figure 6 and present the posterior distributions of key retrieved parameters in Figure 7.666The best-fit model spectra and the fitted T-P profiles of all individual retrieval runs, as well as the QEQ-1 parameter posteriors, are accessible through this Zenodo repository, https://doi.org/10.5281/zenodo.15654303
@@ 2. $\bullet$
The thick clouds and potential high-altitude gray cloud content of 2MASS 1207 b, as suggested by our retrievals, may be contributed by this object’s ongoing accretion (Luhman et al., 2023; Marleau et al., 2024; Manjavacas et al., 2024) of dust and gas from the circumstellar disk of 2MASS 1207 A (e.g., Ricci et al., 2017) which has an estimated dust mass of $\sim 0.1$ M (e.g., Mohanty et al., 2013; Ricci et al., 2017). Additionally, 2MASS 1207 b may host its own disk, though if present, it likely has a low dust mass below 1 Mmoon (Ricci et al., 2017). Similar to events that have influenced Jupiter — such as the impacts of Comet Shoemaker-Levy 9 (e.g., Hammel et al., 1995) — such processes may also influence the atmospheric evolution of 2MASS 1207 b. Follow-up studies similar to Madurowicz et al. (2023) would be valuable in assessing whether accretion activity can match the observed cloud properties of 2MASS 1207 b, as well as the present-day atmospheric compositions of both 2MASS 1207 A and b. Furthermore, future retrieval studies that incorporate both NIRSpec and MIRI spectra of 2MASS 1207 b will be crucial for more precisely constraining the type and amount of silicate clouds in its atmosphere and also the presence and properties of this object’s circumsecondary disk, which, if present, could contribute to the flux at MIRI wavelengths (Luhman et al., 2023).
@@ 7.2.2 Absent CH4 absorption due to the hot atmospheric thermal structure
In Figure 10, we compare the retrieved T-P profile with a critical curve corresponding to an equilibrium situation where the volume mixing ratios of CH4 and CO are equal. This curve, $\log{(P)}=5.05-5807.5/T+0.5\times$[M/H], was derived by Fortney et al. (2020). Deviations from this curve lead to either a decrease in CH4 abundance at higher temperatures and lower pressures, or a decrease in CO abundance at lower temperatures and higher pressures. Objects with thermal structures hotter than this CO–CH4 transition curve tend to exhibit weak or absent 3.3 $\mu$m CH4 absorption, while cooler objects display stronger CH4 absorption. We demonstrate this effect in Figures 1112 using the Sonora Diamondback atmospheric models (Morley et al., 2024).
@@ 7.5 Atmospheric Composition
Focusing on retrieval runs with the seven lowest BIC values (QEQ-1 to QEQ-7), the inferred [M/H] of 2MASS 1207 b ranges from $-0.12$ to $+0.10$ dex, and C/O from 0.44 to 0.52 (Table3). These retrievals incorporate models of atmospheric inhomogeneity, such as patchy clouds or a combination of patchy clouds and hot spots (Figure 5), and explore various assumptions about cloud composition and parameter priors. Despite these differences in model assumptions, their consistent results support the robustness of the near-solar [M/H] and C/O values. Specifically, the most favored retrieval setup, QEQ-1, yields a [M/H]$=-0.05\pm 0.03$ dex and a C/O of 0.440±0.012plus-or-minus0.4400.0120.440$\pm$ 0.0120.440 ± 0.012.999The inferred atmospheric compositions of 2MASS 1207 b — and more broadly, those of exoplanets and brown dwarfs — are derived from disk-integrated spectra, which average out potential three-dimensional spatial variations in atmospheric abundances. Studies of Jupiter suggested that certain gas abundances retrieved from disk-integrated spectra can differ from those of localized regions (e.g., Gapp et al., 2024; Irwin et al., 2025). Comparing the globally and locally inferred atmospheric properties of solar-system planets, such as Jupiter, provides crucial context for interpreting exoplanet and brown dwarf atmospheres and remains a key topic of ongoing research. As a member of the TW Hydrae association, 2MASS 1207 b’s near-solar [M/H] is consistent with stellar members of other nearby young moving groups (e.g., Santos et al., 2008; Ammler-von Eiff & Guenther, 2009; Viana Almeida et al., 2009; D’Orazi et al., 2011; Biazzo et al., 2011, 2012a, 2017).